OpenStudy (sleepyhead314):

x f(x) f'(x) 0 2 5 4 -3 11 Consider the twice-differentiable function \(f\) that has values of \(f\) and its derivative for x=0 and x=4 listed in the table above. \[If~~\int\limits_{0}^{4} f(x)~dx = 8\]what is the value of \[\int\limits_{0}^{4} x~f'(x)~dx~~~?\]

OpenStudy (sleepyhead314):

a) -20 b) -13 c) -12 d) -7 e) 36

OpenStudy (ikram002p):

do integration bye parts :- \(\int\limits_{0}^{4} x~f'(x)~dx = xf(x)|_0^4 - \int_0^4 f(x) dx\)

OpenStudy (sleepyhead314):

This is what I've tried so far... integration by parts ∫xf'(x) dx u = x v = f(x) du = dx dv = f'(x) dx uv - ∫v du (x)(f(x)) - ∫f'(x) dx 4(-3) - 8 -12 - 8 -20 ?

OpenStudy (ikram002p):

ur correct :)

OpenStudy (sleepyhead314):

Thank you! :D

OpenStudy (ikram002p):

but be carfull in the integral its f(x) not f'(x ) (x)(f(x)) - ∫f(x) dx

OpenStudy (ikram002p):

np :)

OpenStudy (sleepyhead314):

ah right typo :P

OpenStudy (ikram002p):

it should be like this :- ( for net time ) u = x dv = f(x) dx du = dx v = f(x)

OpenStudy (ikram002p):

next*

OpenStudy (sleepyhead314):

alright, I'll keep that in mind ^_^

OpenStudy (ikram002p):

^_^